OpenStudy (anonymous):

What value of n solves the equation? 3n = 1 over 81 n =

OpenStudy (lxelle):

Divide 3 on both sides.

OpenStudy (lxelle):

to get only n.

OpenStudy (lxelle):

caz \[\frac{ 3n }{ 3 } = n\]

OpenStudy (phi):

you should put in an ^ to show it's an exponent 3^n = 1/81

OpenStudy (anonymous):

this isnt making any sense to me

OpenStudy (phi):

the first "trick" is factor 81. can you do that ?

OpenStudy (anonymous):

like get the factors of 81?

OpenStudy (phi):

yes, the prime factors

OpenStudy (anonymous):

1, 3, 9, 27, 81

OpenStudy (phi):

I meant what prime numbers do you multiply together to get 81 ? for 6, it would be 2*3 for 8 it would be 2*2*2 etc

OpenStudy (anonymous):

oh, okay... but how is this in relation to my question? /.\ Sorry imn really confused rn

OpenStudy (phi):

you will see in a minute

OpenStudy (phi):

you find them by dividing 81 by the first prime that works: by 3 3*27 then simplify 27, to get 3*3*9 and finally 3*3*3*3

OpenStudy (phi):

now write 3*3*3*3 in "short-hand" using exponents can you do that ?

OpenStudy (anonymous):

what do you mean by in short hand?

OpenStudy (phi):

using exponents

OpenStudy (phi):

people invented exponents to make it easy to write 2*2*2*2 as 2^4

OpenStudy (anonymous):

oh okay. Yes... so like 2 to the 4 power. 2 tot eh 4th power is 16... sorry I woke up like 10 minutes ago... and im still tired... oh god cx

OpenStudy (phi):

2 was an example so far we know 81 = 3*3*3*3 use exponents to re-write the right-hand side

OpenStudy (anonymous):

yes... but how am I supposed to re-write 1 over 81?

OpenStudy (phi):

one step at a time. can you write 3*3*3*3 using exponents?

OpenStudy (anonymous):



OpenStudy (phi):

ok, but you can type it as 3^4 in your problem, we can write \[ 3^n = \frac{1}{81} \\ 3^n = \frac{1}{3^4} \] ok ?

OpenStudy (anonymous):

oh, okay

OpenStudy (anonymous):



OpenStudy (phi):

the problem is as I posted \[ 3^n = \frac{1}{3^4} \] However, there is a rule using exponents. you can "flip" a fraction and make the exponent negative. Like this: \[ 3^n = 3^{-4} \]

OpenStudy (phi):

now you "match up" 3^n with 3^(-4) the "n" matches with the -4